Find the minimum sample size needed to be 95% confident that the sample's variance is within 50% of the population's variance.
Answer #1

Z value for 95% confidence is 1.96

E = margin of error = 0.5
Required sample size to estimate the variance:

n = (Z*S/E)^2 = (1.96 * 1 /0.5)^2 = 16

Hope this will be helpful. Thanks and God Bless You:)